1. What is the coefficient of performance of an ideal heat pump that has heat transfer from a cold temperature of −25.0ºC to a hot temperature of 40.0ºC?
  2. Suppose you have an ideal refrigerator that cools an environment at−20.0ºC and has heat transfer to another environment at 50.0ºC. What is its coefficient of performance?
1
Expert's answer
2021-07-05T17:59:45-0400

1)


"\frac{1}{COP}=1-\frac{273-25}{273+40}\\COP=4.82"

2)


"\frac{1}{COP+1}=1-\frac{273-20}{273+50}\\COP=3.61"